### 11 Jan 2024 v1

# LogPTR: Variable-Aware Log Parsing with Pointer Network

Abstract
Due to the sheer size of software logs, developers rely on automated log analysis. Log parsing, which parses semi-structured logs into a structured format, is a prerequisite of automated log analysis. However, existing log parsers are unsatisfactory when applied in practice because: 1) they ignore categories of variables, and 2) have poor generalization ability. To address the limitations of existing approaches, we propose LogPTR, the first end-to-end variable-aware log parser that can extract the static and dynamic parts in logs, and further identify the categories of variables. The key of LogPTR is using pointer network to copy words from the log message. We have performed extensive experiments on 16 public log datasets and the results show that LogPTR outperforms state-of-the-art log parsers both on general log parsing that extracts the log template and variable-aware log parsing that further identifies the category of variables.
Index Terms—  software logs, log parsing, pointer network, deep learning
1 Introduction
Logs play an important role in software systems to record system execution behaviors. By inspecting system logs, engineers can detect anomalies [1], localize software bugs [2], or troubleshoot problems [3] in the system. However, facing the rapid growth volume of logs, it is becoming more and more challenging to identify valuable information from the enormous log data, even for experienced engineers. To tackle this problem, automated log analysis has emerged in recent years, aiming to automatically analyze the log data [4].
The fundamental step for automated log analysis is log parsing, which parses the semi-structured log messages into a structured format [4]. Log messages are generated from logging statements in the source code. A log message usually contains the log header (e.g., time, levels) and the message content, which further consists of two parts: 1) static log template describing system events; 2) dynamic variables, which vary during runtime and reflect system runtime status. The goal of log parsing is to extract the static log template and dynamic log variables from a log message, as shown in Figure 1.
Fig. 1: An example of log parsing from Spark.
Fig. 2: The model architecture of LogPTR.
To achieve automated log parsing, many approaches have been proposed using different techniques [4], such as similarity-based clustering (e.g., LKE [5] and LenMa [6]), frequent pattern mining (e.g., LFA [7] and Logram [8]), heuristics-based searching (e.g., Spell [9], Drain [10] and Brain [11]), and deep learning-based parsing (e.g., Uniparser [12], LogPPT [13], and Semlog [14]). Although making progress, existing log parsers are still unsatisfactory when applied in practice. We have identified the following limitations of the existing log parsers:
  • Ignoring categories of variables. The categories of variables record valuable information to assist log understanding and analysis [15]. However, the goal of most log parsers is to extract all dynamic variables of logs and output the remaining static words, which ignores the categories of variables.
  • Poor generalization ability. Most log parsers require handcrafted rules for data pre-processing (e.g., defining regular expressions) and hyperparameter settings (e.g., the number of clusters or similarity threshold). When applying to a new log dataset, due to different logging formats and behaviors, the time-consuming adjustment of handcrafted rules and hyperparameters makes it suffer from system migrations.
To tackle the aforementioned limitations, we propose LogPTR, the first end-to-end variable-aware log parser. LogPTR is able to identify which words are static constants, which words are variables and what are their corresponding categories. First, we add label tokens at the beginning of log messages to indicate the categories of variables. Then, LogPTR can generate the variable-aware log template by copying words from the log message with the pointer mechanism. LogPTR does not require any handcrafted rules and uses the same set of hyperparameter values for each dataset, which is robust across different types of logs, and more generalized than existing approaches.
We have evaluated LogPTR on 16 public log datasets [16]. LogPTR achieves an average group accuracy of 0.989 and parsing accuracy of 0.972 for general log parsing, which outperforms state-of-the-art log parsers. In addition, LogPTR can efficiently identify the categories of variables in logs to perform variable-aware log parsing and achieves an average parsing accuracy of 0.972, which is higher than the baseline. Moreover, LogPTR is also robust across different log datasets.
To summarise, our main contributions are as follows:
  • We propose LogPTR, a variable-aware log parser that can identify log templates, variables, and categories of variables. LogPTR does not require any handcrafted rules and uses the same set of hyper-parameter values for every dataset, thus can quickly adapt to new log datasets. To the best of our knowledge, LogPTR is the first end-to-end log parser.
  • We evaluate LogPTR on 16 public log datasets and the results show that LogPTR outperforms prior state-of-the-art approaches both in general log parsing and variable-aware log parsing.
2 METHODOLOGY
2.1 Problem Definition
Intuitively, log messages can be regarded as a special type of natural language due to their semi-structured essence of mixing unstructured natural language and structured variables. Thus, we formulate log parsing as a text summarization problem, where the output vocabulary consists of the tokens in the log message and the label tokens indicating categories of variables. Specifically, for general log parsing, we only use one label token to indicate variables, i.e., ignore categories of variables. While for variable-aware log parsing, we add ten label tokens to indicate the categories of variables [15].
Formally, suppose there are m categories of variables, denoted as $C=\{c_{1},...,c_{m}\}$ and a raw log message consists of n tokens after tokenization, denoted as $T=\{t_{1},..,t_{n}\}$. Our proposed log parser takes $X=[C,T]$ as input and generates the output $Y=\{y_{1},...,y_{k}\}\ y_{i}\in[1,m+n]$, where $y_{i}$ indicates the index of the input. Based on the indices, we can output the variable-aware log template $L=\{X{y_{1}},...,X{y_{k}}\}$ as the final result.
Table 1: Accuracy comparison with the state-of-the-art log parsers on general log parsing.
Dataset AEL LenMa Spell Drain Logram LogPPT VALB* LogPTR
GA PA GA PA GA PA GA PA GA PA GA PA PA GA PA
Android 0.642 0.559 0.778 0.722 0.891 0.241 0.765 0.730 0.677 0.428 0.885 0.767 0.935 0.968 0.948
Apache 0.984 0.987 0.984 0.293 0.301 0.285 1 1 0.297 0.509 1 0.994 1 1 1
BGL 0.491 0.410 0.316 0.154 0.85 0.329 0.955 0.444 0.218 0.170 0.954 0.97 0.913 0.903 0.937
Hadoop 0.677 0.422 0.667 0.242 0.533 0.196 0.948 0.439 0.283 0.370 0.994 0.895 0.977 0.995 0.965
HDFS 0.626 0.630 0.998 0.125 1 0.487 0.998 0.959 0.012 0.018 1 0.902 0.970 0.997 0.996
HealthApp 0.57 0.175 0.166 0.289 0.961 0.152 0.644 0.241 0.258 0.263 1 0.789 0.993 1 0.978
HPC 0.731 0.698 0.681 0.671 0.657 0.532 0.741 0.672 0.742 0.679 0.943 0.947 0.992 0.994 0.979
Linux 0.404 0.239 0.238 0.132 0.162 0.109 0.422 0.194 0.201 0.185 0.934 0.949 0.965 1 0.968
Mac 0.583 0.223 0.648 0.155 0.724 0.033 0.711 0.277 0.551 0.252 0.78 0.673 0.866 0.982 0.897
OpenSSH 0.198 0.421 0.927 0.155 0.488 0.127 0.996 0.534 0.343 0.482 0.628 0.976 0.982 0.997 0.989
OpenStack 0.266 0.187 0.213 0.191 0.245 0.000 0.224 0.187 0.241 0.112 0.989 0.907 0.938 1 0.994
Proxifier 0.495 0.506 0.495 0.506 0.527 0.478 0.527 0.527 0.016 0.000 1 1.000 1 1 1
Spark 0.415 0.381 0.869 0.023 0.92 0.336 0.905 0.376 0.282 0.275 0.999 0.991 0.993 1 0.992
Thunderbird 0.65 0.203 0.943 0.171 0.856 0.039 0.96 0.191 0.129 0.128 0.679 0.926 0.881 0.986 0.939
Windows 0.685 0.389 0.287 0.266 0.99 0.004 0.994 0.696 0.694 0.374 0.991 0.983 0.992 0.998 0.988
Zookeeper 0.657 0.527 0.894 0.457 0.987 0.453 0.967 0.498 0.724 0.516 0.994 0.99 0.983 1 0.989
Average 0.567 0.435 0.631 0.284 0.693 0.237 0.797 0.498 0.354 0.297 0.923 0.916 0.961 0.989 0.972
    As open-source code is unavailable, we use the paper’s results, which evaluated VALB using only the PA metric.
2.2 Model Architecture
LogPTR adopts sequence-to-sequence architecture. The encoder consists of an embedding layer and a bidirectional LSTM (Bi-LSTM) layer, while the decoder consists of a uni-directional LSTM with the same hidden size as that of the encoder, and a pointer mechanism to copy words from the log message. The architecture of LogPTR is shown in Figure 2.
Embedding Layer. Engineers can create an infinite variety of variable names or abbreviations, which are far beyond the scale of common words. In addition, many variables in logs are digits (e.g., “23869471”) or special symbols (e.g., “/etc/data/”), which can almost have unlimited potential of creating “new words” based on different system behaviors. This may incur the out-of-vocabulary (OOV) problem when encoding logs [12, 15]. To solve the problem, we use WordPiece [17] to tokenize the log message into a sequence of subwords. Then, the embedding layer encodes each subword into a high-dimensional embedding vector as the next layer input.
Bi-LSTM Layer. Similar to sentences in natural language, the log is a series of words that have sequential dependencies between the words. Hence, we use a Bi-LSTM to capture the dependencies between words in logs. Specifically, the subword vector is fed into the Bi-LSTM layer. Afterward, the hidden state $e_{i}$ of Bi-LSTM is used for the pointer mechanism to generate words.
Pointer Mechanism. We use the pointer mechanism [18] to generate the variable-aware log template by creating pointers to words in the input sequence. Specifically, on each step $t$, the decoder receives the word embedding of the previous word and has decoder hidden state $d_{t}$. We obtain the following output distribution over the vocabulary of inputs:
$\small u^{t}_{i}$ $=v^{T}tanh(W_{1}e_{i}+W_{2}d_{t})$ (1)
$P(y_{t}|y_{<t},X)$ $=softmax(u^{t})$ (2)
where $v$, $W_{1}$, and $W_{2}$ are learnable parameters.
During training, LogPTR is trained with the maximum likelihood objective, defined as:
$\small\mathcal{L}=-{\textstyle\sum_{t=0}^{T}}logP(y_{t}|y_{<t},X)$ (3)
3 EXPERIMENTAL DESIGN
3.1 Datasets
Our experiments are conducted on 16 public log datasets from LogHub [16], the most widely-used benchmark in log analysis. Specifically, we use the corrected version of the 16 datasets from [19] to evaluate general log parsing. Furthermore, we use the dataset from [15] to evaluate variable-aware log parsing, where the 16 datasets in LogHub were manually annotated including 10 categories of variables. For each dataset, we randomly split it into training (20%), validation (20%), and testing data sets (60%).
3.2 Evaluation Metrics
Following recent studies [12, 13, 19, 20], we apply two widely used metrics in our evaluation, including:
Group Accuracy (GA): Group Accuracy [20] is the most commonly used metric for log parsing. GA considers log parsing as a clustering process where log messages with different log templates are clustered into different groups [19]. GA is defined as the ratio of “correctly parsed” log messages over the total number of log messages, where a log message is considered “correctly parsed” if and only if it is grouped with other log messages consistent with the ground truth.
Parsing Accuracy (PA): Parsing Accuracy (or Message-Level Accuracy [12]) is defined as the ratio of “correctly parsed” log messages over the total number of log messages, where a log message is considered to be “correctly parsed” if and only if every token of the log message is correctly identified as a template or variable in general log parsing, or further all the categories of variables are also correctly identified in variable-aware log parsing.
3.3 Implementation Details
We conduct experiments on a GPU Server equipped with NVIDIA Tesla V100 GPU and CUDA 11.7. We implement LogPTR with Python 3.8 and PyTorch 2.0. For the embedding layer, we set the dimension of word embedding as 256. For the LSTM layer, we set the hidden size to 256. To mitigate the impact of overfitting, we apply the dropout method for the embedding layer and LSTM layer, with a dropout rate of 0.2. We utilize Adam optimizer with an initial learning rate of 0.001. During the training phase, we set the batch size as 32 and train the model for 100 epochs.
4 EXPERIMENTAL RESULTS
4.1 Effectiveness on General Log Parsing
Accuracy. To use LogPTR as a general log parser that identifies all the variables, we set the label token as [VAR] to indicate the token is a variable, regardless of its categories. We compare the accuracy of LogPTR with seven state-of-the-art log parsers (including AEL [21], LenMa [6], Spell [9], Drain [10], Logram [8], LogPPT [13], and VALB [15]) on all 16 log datasets. These log parsers cover most types of parsing techniques and achieve good results on public datasets [19, 20]. Table 1 presents the accuracy of LogPTR and the other state-of-the-art log parsers on general log parsing.
From the results, we can see that LogPTR outperforms the baselines on almost all datasets in the two metrics. Specifically, in terms of GA, LogPTR achieves the highest average accuracy of 0.989 and the best results on 14 out of 16 datasets. It is worth noting that LogPTR achieves an accuracy of over 0.95 on 15 datasets and achieves 1.0 accuracy on 7 datasets among them, which is significantly superior to existing log parsers. In terms of PA, LogPTR also achieves the highest average accuracy of 0.972. Overall, the experimental results confirm that LogPTR is effective in general log parsing which groups logs into the same templates and identifies correct templates and variables.
Fig. 3: Robustness comparison with the state-of-the-art log parsers across different log datasets.
Robustness. LogPTR explicitly aims at supporting a broad range of diverse log datasets because robustness is crucial to the use of a log parser in practice [20]. Therefore, we analyze and compare the robustness against different types of logs of LogPTR with that of the baselines, as shown in Figure 3.
From the results, we can see that LogPTR outperforms the baselines in terms of robustness across different log types. Existing methods require different regular expressions for pre-processing and different hyper-parameter values, which involve domain-specific knowledge and labor-intensive model tuning. Thus, they perform inconsistently on different datasets. In contrast, LogPTR does not require manually defining regular expressions and achieves the smallest variance over different datasets. Moreover, LogPTR uses the same set of hyper-parameter values in the training phase and does not require re-adjustment for each dataset. Overall, the experimental results confirm that LogPTR is robust and can be applied to different log datasets with low effort.
Table 2: Accuracy comparison with the state-of-the-art log parsers on variable-aware log parsing.
Dataset VALB LogPTR
Android
0.916
0.943
Apache
0.993
1
BGL
0.896
0.959
Hadoop
0.968
0.968
HDFS
0.965
0.998
HealthApp
0.988
0.971
HPC
0.99
0.992
Linux
0.959
0.961
Mac
0.862
0.866
OpenSSH
0.976
0.994
OpenStack
0.932
0.998
Proxifier
1
1
Spark
0.991
0.997
Thunderbird
0.878
0.925
Windows
0.99
0.983
Zookeeper
0.981
0.989
Average
0.955
0.972
4.2 Effectiveness on Variable-aware Log Parsing
Apart from identifying static and dynamic parts in logs (i.e., general log parsing), LogPTR can also identify the categories of variables (i.e., variable-aware log parsing). In this RQ, we compare the accuracy of LogPTR on variable-aware log parsing with VALB [15], which is the only baseline that can identify the categories of variables. Table 2 presents the accuracy of LogPTR and VALB on variable-aware log parsing.
From the results, we can see that LogPTR achieves a high parsing accuracy ranging from 0.866 in Mac to 1 in Apache, which is also close to the parsing accuracy as discussed in RQ1. The average parsing accuracy of LogPTR is 0.972, which is higher than the baseline (i.e., 0.955). Moreover, LogPTR achieves the best results on 14 out of 16 datasets. As for the robustness, we calculate the standard deviation (STD) of parsing accuracy, and LogPTR achieves a smaller STD than the baseline (0.036 versus 0.045). Overall, the experimental results confirm that apart from general log parsing, LogPTR can also efficiently identify the categories of variables in logs to perform variable-aware log parsing.
5 CONCLUSION
This paper proposes LogPTR, the first end-to-end variable-aware log parser that can extract static and dynamic parts in logs, and further identify categories of variables. The key of LogPTR is using pointer network to copy words from the log message. Experimental results on 16 public log datasets show that LogPTR outperforms the state-of-the-art log parsers both on general log parsing and variable-aware log parsing.